countryboy:

What is the approximate volume of a cone with a height of 6 mm and radius of 18 mm? Use 3.14 to approximate pi, and express your final answer to the nearest hundredth. Enter your answer in the box. mm³

ThisGirlPretty:

@Shadow Whenever you get the chance sir >~<

iosangel:

\[V=πr2h/3\] \[3.14∗182∗6/3\] =3.14*324*6/3 =1017.36*6/3 =1017.36*2 =2034.72